OpenStudy (anonymous):

What is the area of the parallelogram with sides of 5" and 9"? It comes with this pic...

OpenStudy (anonymous):

I don't get how it corresponds to the question...

OpenStudy (freethinker):

do you know the formula to obtain the area of a parallelogram? even if you didn't you should be able to derive from two known areas - rectangle and right triangle area of rectangle = length x width are of triangle 1/2(length of the base x height), which in your case two triangles.

OpenStudy (anonymous):

Yeah but how does the sides of 5" and 9" in the question correspond to the values shown in the pic?

OpenStudy (freethinker):

you are given the dimensions 10, 8 and 20 this implies that your rectangle would have the dimensions of 8 for the width and 20 for the length. if you know a bit more about pythagorean theorem, you would be able to find the length of the base of the triangle.

OpenStudy (freethinker):

the square root of the sum of the squares of short sides of a right triangle gives you the length of the hypotenuse. does this sound familiar to you? \[c \sqrt{a^2+b^2}\]

OpenStudy (anonymous):

is that the same as a^2+b^2=c^2?

OpenStudy (freethinker):

this question is mainly testing your ability to put the knowledge you've got to figure out the area of a parallelogram without giving you directly the dimensions of a parallelogram.

OpenStudy (freethinker):

yes that is the same as the equation you provided. the reason you have to obtain the square root of c is because having it in squared form is simply not the value of c in a right-triangle.

OpenStudy (anonymous):

hmm okay so using the values given, the missing length of the triangle is 6 isn't it? 6-8-10 Pythagorean triple?

OpenStudy (freethinker):

and you remember one of the rules in solving math is that whatever you do on one side of the equation, you must also on the other. this is to maintain the balance and so that you don't change the equation itself.

OpenStudy (freethinker):

yes, because of the pythagorean triple, you are able to easily obtain the value of one of the dimensions.

OpenStudy (anonymous):

so is the base of the parallelogram 26 now?

OpenStudy (freethinker):

exactly. the length of the longer parallel sides are both 26

OpenStudy (anonymous):

so given that the area of a parallelogram is bxh, 26x8=208u^2 is the answer?

OpenStudy (freethinker):

yes, whatever unit is squared when calculating for an area.

OpenStudy (freethinker):

by the way you posted this question in the wrong category. This is biology section, not math…

OpenStudy (anonymous):

yes I just saw that, my bad! and thank you so much for your help, I appreciate it!

OpenStudy (goformit100):

@Preetha SEE it's a Mathemaical question posted in the wrong section

OpenStudy (umarsback):

Goformit has been acting weird since he's went green...

OpenStudy (goformit100):

Sir I am only following http://openstudy.com/code-of-conduct.